You will need to use the formula 4.9t^2 + v0t = s, where t is the time in seconds, v0 is the initial velocity, and s is the distance in meters.
Approximately how far will an object fall in 13 seconds (t) when thrown downward at an initial velocity of 10 meters per second (v0) from a plane?

Found 2 solutions by stanbon, Alan3354:
Answer by stanbon(75887)

You will need to use the formula 4.9t^2 + v0t = s(t), where t is the time in seconds, v0 is the initial velocity, and s is the distance in meters.
Approximately how far will an object fall in 13 seconds (t) when thrown downward at an initial velocity of 10 meters per second (v0) from a plane?
------
s(13) = 4.9(13)^2 + 10(13)
S(13) = 958.1 ft.
=====================
Cheers,
Stan H.

Answer by Alan3354(69443)

Approximately how far will an object fall in 13 seconds (t) when thrown downward at an initial velocity of 10 meters per second (v0) from a plane?
------
s(13) = 4.9(13)^2 + 10(13)
S(13) = 958.1 meters